# Classifying locally distinguishable sets: No activation across bipartitions

### v1

@@ IV Non-activable entangled states
Here we consider several sets that are local and non-activable by LOCC, i.e., $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{j})=0$, for different sets $\mathcal{S}_{j}$. $|^{\theta}\mathcal{W}_{ij,kl}^{\pm}\rangle_{AB}=|\theta\rangle_{A}|{\mathcal{X}}_{ij}^{\pm}\rangle_{B}+|\bar{(\theta+2)}\rangle_{A}|{\mathcal{X}}_{kl}^{\pm}\rangle_{B},\;$ and $|^{\theta}\mathcal{W}_{ij,m}^{\pm}\rangle_{AB}=|\theta\rangle_{A}|{\mathcal{X}}_{ij}^{\pm}\rangle_{B}+|\bar{(\theta+2)}\rangle_{A}|{\mathcal{X}}_{m}\rangle_{B},\;$ also $|^{\theta}\mathcal{\bar{W}}_{ij,kl}^{\pm}\rangle_{AB}=|{\mathcal{\xi}}_{ij}^{\pm}\rangle_{A}|\theta\rangle_{B}+|{\mathcal{\xi}}_{kl}^{\pm}\rangle_{A}|\bar{(\theta+2)}\rangle_{B},\;$ and $|^{\theta}\mathcal{\bar{W}}_{ij,m}^{\pm}\rangle_{AB}=|{\mathcal{\xi}}_{ij}^{\pm}\rangle_{A}|\theta\rangle_{B}+|{\mathcal{\xi}}_{m}\rangle_{A}|\bar{(\theta+2)}\rangle_{B},\;$, for $\theta=0,1$ and $\bar{(\theta+2)}$ denotes $(\theta+2)$ modulo d. Here we consider the set $\mathcal{S}_{5}=\left\{\left|\phi_{i}\right\rangle_{AB}\right\}\in\mathbf{\mathbf{C}^{4}\otimes\mathbf{C}^{4}}$, which contains product states as well as entangled states. The set is given by-
@@
The set $\mathcal{S}_{5}$ does not possess any activable nonlocality under orthogonality-preserving LOCC. That is, its hidden nonlocality $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{5})=0.$
@@
Here the states of (4) are nothing but the superposition of states of (1). The only difference between them is that one contains entangled states and the other contains only product states. So, the outline of the proof is similar to that of the Proposition 1. Without loss of generality, let us assume that Alice goes first. Considering the states $|0\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B},\;|1\rangle_{A}|{\mathcal{X}}_{3}\rangle_{B},$ we have $\left\langle 0\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|1\right\rangle_{A}\left\langle 2+3|3\right\rangle_{B}=0,$ which implies that , $m^{a}_{01}=m^{a}_{10}=0.$
In the same way, for the states $|^{0}\mathcal{W}_{01,23}^{+}\rangle_{AB},\;|0\rangle_{A}|{\mathcal{X}}_{23}^{\pm}\rangle_{B}$ and $|^{1}\mathcal{W}_{12,3}^{+}\rangle_{AB}$, we have $m^{a}_{02}=m^{a}_{20}=0,m^{a}_{03}=m^{a}_{30}=0$, respectively. Similarly, by choosing appropriate pair of states we get $m^{a}_{ij}=m^{a}_{ji}=0,\forall i\neq j$. Therefore, ${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}$ is diagonal and ${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}$ $=\operatorname{diag}\left(\delta_{0},\delta_{1},\delta_{2},\delta_{3}\right).$
@@
Figure 4: Tile structure of the states in ${\mathbf{C}}^{4}\otimes{\mathbf{C}}^{4}$, given in (4). The indices of the tiles comes from the order the states of $\mathcal{S}_{5}$ consecutively. The color of the tiles represents the possibility of measurement setup for both the parties.
Next considering $|^{0}\mathcal{\bar{W}}_{12,3}^{+}\rangle_{AB}$ and $|{\mathcal{\xi}}_{12}^{-}\rangle_{A}|0\rangle_{B},$ we get $\left\langle 1+2\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|1-2\right\rangle_{A}\langle 0|0\rangle_{B}=0,$ i.e., $\left\langle 1\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|1\right\rangle-\left\langle 2\left|{\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}\right|2\right\rangle=0$. Thus, $m^{a}_{11}=m^{a}_{22}$. By using the states $|{\mathcal{\xi}}_{23}^{\pm}\rangle_{A}|1\rangle_{B},$ we finally get $m^{a}_{11}=m^{a}_{22}=m^{a}_{33}.$ Therefore, ${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}$ $=\operatorname{diag}\left(\delta_{0},\gamma,\gamma,\gamma\right).$ If possible, let us assume that $\delta_{0}\neq 0$ and $\gamma\neq 0$. Then after Alice’s measurement, Bob should do a nontrivial operation on his own subsystem according to Alice’s result. We denote $\mathcal{M}_{B}^{m}\;(where,\;{\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}=(m^{b}_{ij})_{4\times 4})$ as Bob’s operator. As we have discussed above, by choosing suitable pair of states we can conclude that all the off-diagonal elements of ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}$ are equal to 00. Similarly, for the diagonal elements as we have discussed above, if we consider the states $|^{0}\mathcal{W}_{01,23}^{\pm}\rangle_{AB},\;|0\rangle_{A}|{\mathcal{X}}_{23}^{\pm}\rangle_{B},\;|{\mathcal{\xi}}_{3}\rangle_{A}|0\rangle_{A}$ and $|1\rangle_{A}|{\mathcal{X}}_{12}^{-}\rangle_{B}$, we finally get, $m^{b}_{00}=m^{b}_{11}=m^{b}_{22}=m^{b}_{33}.$ Therefore, ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}$ is proportional to the identity operator, i.e., ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}=\lambda_{0}I$, which is the trivial operator and this contradicts our assumption that $\mathcal{S}_{5}$ is initially local. So, either $\delta_{0}=0$ or $\gamma=0.$ Notice that this result also suggests us that these states cannot be distinguished locally if Bob goes first. Now it is clear that if Alice goes first with a diagonal operator, i.e., $\delta_{0}=\gamma=1$, then the above set of states cannot be distinguished. So, Alice has to do non-trivial measurement first and this only happens when any one of $\delta_{0}$, $\gamma$ is not equal to zero. For that Alice only has two outcome measurement operators: $\quad{\mathcal{M}_{A}^{1}}^{\dagger}\mathcal{M}_{A}^{1}=\operatorname{diag}(1,0,\ldots,0)$ and ${\mathcal{M}_{A}^{2}}^{\dagger}\mathcal{M}_{A}^{2}$ $=\operatorname{diag}(0,1,1\ldots,1)$, see Fig. 4. If the outcome ‘$1$’ click, Bob is able to distinguish the remaining states by projecting onto $\left|0\pm 1\right\rangle$ and $\left|2\pm 3\right\rangle$. If the measurement outcome is ‘$2$’, it will isolate the remaining states. It is then Bob’s turn to do measurement. Following the method we used above, we can similarly prove that Bob’s measurement must be ${\mathcal{M}_{B}^{1}}^{\dagger}\mathcal{M}_{B}^{1}=\operatorname{diag}(1,0,0,0)$ and ${\mathcal{M}_{B}^{2}}^{\dagger}\mathcal{M}_{B}^{2}$ $=\operatorname{diag}(0,1,1,1)$. The process will repeat a finite number of times, and for each measurement outcome for both parties, the set $\mathcal{S}_{3}$ transforms only to a distinguishable set. This implies the fact that, if the set is distinguishable (local), then for all possible nontrivial measurements, it is impossible to transform the set into an indistinguishable one. That is in other words, the set $\mathcal{S}_{5}$ is not activable through orthogonality-preserving LOCC. This completes the proof. ∎
@@
Figure 5: Tiles representation of states in $\mathbf{\mathbf{C}^{6}\otimes\mathbf{C}^{6}}$, indexed by states of $\mathcal{S}_{6}$ in order and colored according to the possibility of simultaneous local measurements by both parties.
@@
It is not very difficult to construct the set $\mathcal{S}_{5}$ in arbitrary higher dimensions from its hereditary symmetry. For the case of higher dimensions, the only change will be the number of classical rounds required for discrimination task. Eventually for higher dimensions, the LOCC round numbers drastically increase for the corresponding tasks, but for each round the post-measurement states becomes distinguishable (local). Also, one can find the trade off between the dimensions of the systems and the corresponding required LOCC round number for the discrimination tasks. Now, consider the set $\mathcal{S}_{6}=\left\{\left|\phi_{i}\right\rangle_{AB}\right\}\in\mathbf{\mathbf{C}^{6}\otimes\mathbf{C}^{6}}$, by
@@
By the similar technique as given for (4), it is possible to show that the set $\mathcal{S}_{6}$ does not possess any activable nonlocality under orthogonality-preserving LOCC, i.e., $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{6})=0$. See Fig. 5. Here we now discuss the following hierarchy. The sets of states considered in (5) and (3) are equally local when considered with respect to perfect discrimination by LOCC, as in both cases, the sets are perfectly distinguishable by LOCC. But the consideration of hidden nonlocality provides us the privilege to put a hierarchy among the sets. Precisely, we can claim that the sets of (5) are more local compared to those of (3), because the latter class contains hidden nonlocality while for the former case there is no hidden nonlocality though the set of (5) contains entangled states but the set of (3) does not.
@@ V Conclusion
In this manuscript, we have presented structures for locally distinguishable product states and entangled states such that they cannot be transformed to a locally indistinguishable set under orthogonality-preserving LOCC. Furthermore, we have constructed sets of multipartite states which is not activable in any bipartition. In other words, such locally distinguishable sets cannot be transformed to a locally indistinguishable set in any bipartition under orthogonality-preserving LOCC. This is, in fact, the worst case scenario in view of nonlocality activation. We also have constructed locally distinguishable sets which can be transformed to locally indistinguishable sets under orthogonality-preserving LOCC. Then, we have classified the locally distinguishable states by introducing hierarchies. The structures that we have provided here can be easily generalized for high-dimensional Hilbert spaces. Finally, we have compared between locally distinguishable product states and entangled states.
@@ ACKNOWLEDGMENT
SH acknowledges the funding by the European Union under Horizon Europe (grant agreement no. 101080086). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ IV Non-activable entangled states
Here we consider several sets that are local and non-activable by LOCC, i.e., $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{j})=0$, for different sets $\mathcal{S}_{j}$. $\mathbf{|^{\theta}\mathcal{W}_{ij,kl}^{\pm}\rangle_{AB}=|\theta\rangle_{A}|{\mathcal{X}}_{ij}^{\pm}\rangle_{B}+|\bar{(\theta+2)}\rangle_{A}|{\mathcal{X}}_{kl}^{\pm}\rangle_{B}},\;$ and $\mathbf{|^{\theta}\mathcal{W}_{ij,m}^{\pm}\rangle_{AB}=|\theta\rangle_{A}|{\mathcal{X}}_{ij}^{\pm}\rangle_{B}+|\bar{(\theta+2)}\rangle_{A}|{\mathcal{X}}_{m}\rangle_{B}},\;$ also $\mathbf{|^{\theta}\mathcal{\bar{W}}_{ij,kl}^{\pm}\rangle_{AB}=|{\mathcal{\xi}}_{ij}^{\pm}\rangle_{A}|\theta\rangle_{B}+|{\mathcal{\xi}}_{kl}^{\pm}\rangle_{A}|\bar{(\theta+2)}\rangle_{B}},\;$ and $\mathbf{|^{\theta}\mathcal{\bar{W}}_{ij,m}^{\pm}\rangle_{AB}=|{\mathcal{\xi}}_{ij}^{\pm}\rangle_{A}|\theta\rangle_{B}+|{\mathcal{\xi}}_{m}\rangle_{A}|\bar{(\theta+2)}\rangle_{B}},\;$, for $\mathbf{\theta=0,1}$ and $\mathbf{\bar{(\theta+2)}}$ denotes $\mathbf{(\theta+2)}$ modulo $\mathbf{d}$. Here we consider the set $\mathcal{S}_{5}=\left\{\left|\phi_{i}\right\rangle_{AB}\right\}\in\mathbb{C}^{4}\otimes\mathbb{C}^{4}$, which contains product states as well as entangled states. The set is given by-
@@
The set $\mathcal{S}_{5}$ does not possess any activable non-locality under orthogonality-preserving LOCC. That is, its hidden non-locality $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{5})=0.$
@@
The states in (4) can be viewed as superpositions of the product states given in (1), with the key difference being the inclusion of entangled states. Hence, the overall structure of the proof follows similarly to that of Proposition 1, with the key steps sketched below.
Without loss of generality, let us assume that Alice goes first. By analyzing suitable pairs of states, such as $\mathbf{|0\rangle_{A}|{\mathcal{X}}_{23}^{+}\rangle_{B}}$ and $\mathbf{|1\rangle_{A}|{\mathcal{X}}_{3}\rangle_{B}}$, and using the orthogonality-preserving condition, we find that the off-diagonal terms of Alice’s measurement operator must vanish. Repeating this for other carefully chosen state pairs, we conclude that Alice’s operator must be diagonal. Furthermore, symmetry in the state structure leads to equal diagonal entries: $\delta_{1}=\delta_{2}=\delta_{3}=\gamma$. Therefore, we obtain:
${\mathcal{M}_{A}^{m}}^{\dagger}\mathcal{M}_{A}^{m}=\operatorname{diag}(\delta_{0},\gamma,\gamma,\gamma).$
Assuming both $\delta_{0}\neq 0$ and $\gamma\neq 0$, Bob must apply a nontrivial operation. However, a similar analysis on Bob’s side implies that his operator must also be diagonal and further turns out to be proportional to the identity: ${\mathcal{M}_{B}^{m}}^{\dagger}\mathcal{M}_{B}^{m}=\lambda_{0}I,$ which contradicts the requirement of a nontrivial measurement if the set is to be initially locally distinguishable. Hence, either $\delta_{0}=0$ or $\gamma=0$, implying that Alice must begin with a nontrivial two-outcome measurement:
${\mathcal{M}_{A}^{1}}^{\dagger}\mathcal{M}_{A}^{1}=\operatorname{diag}(1,0,0,0),\quad{\mathcal{M}_{A}^{2}}^{\dagger}\mathcal{M}_{A}^{2}=\operatorname{diag}(0,1,1,1).$
If the outcome ‘$1$’ clicks, Bob can distinguish the remaining states by projections onto $\mathbf{\left|0\pm 1\right\rangle}$ and $\mathbf{\left|2\pm 3\right\rangle}$. If the outcome is ‘$2$’, the remaining subspace can again be handled by Bob using a similar measurement structure. See Fig. 4 for the measurement-induced partitioning of the state space.
@@
Figure 4: (Color online) Tile structure of the states in ${\mathbb{C}}^{4}\otimes{\mathbb{C}}^{4}$, given in (4). The indices of the tiles follow the ordering of the states in $\mathcal{S}_{5}$. Tile colors represent the measurement configuration for both parties.
The process continues for a finite number of steps. At each step, the remaining set remains locally distinguishable. Therefore, the set $\mathcal{S}_{5}$ cannot be transformed into an indistinguishable one through any orthogonality-preserving LOCC sequence. This shows that $\mathcal{S}_{5}$ is not activable under LOCC, and this completes the proof. ∎
@@
Figure 5: (Color online) Tiles representation of states in $\mathbb{C}^{6}\otimes\mathbb{C}^{6}$, indexed by states of $\mathcal{S}_{6}$ in order and colored according to the possibility of simultaneous local measurements by both parties.
@@
It is not very difficult to construct the set $\mathcal{S}_{5}$ in arbitrary higher dimensions from its hereditary symmetry. For the case of higher dimensions, the only change will be the number of classical rounds required for discrimination task. Eventually for higher dimensions, the LOCC round numbers drastically increase for the corresponding tasks, but for each round the post-measurement states becomes distinguishable (local). Also, one can find the trade off between the dimensions of the systems and the corresponding required LOCC round number for the discrimination tasks. Now, consider the set $\mathcal{S}_{6}=\left\{\left|\phi_{i}\right\rangle_{AB}\right\}\in\mathbf{\mathbb{C}^{6}\otimes\mathbb{C}^{6}}$, by
@@
By the similar technique as given for (4), it is possible to show that the set $\mathcal{S}_{6}$ does not possess any activable non-locality under orthogonality-preserving LOCC, i.e., $\mathcal{H}^{\text{LOCC}}_{1}(\mathcal{S}_{6})=0$. See Fig. 5. Here we now discuss the following hierarchy. The sets of states considered in (5) and (3) are equally local when considered with respect to perfect discrimination by LOCC, as in both cases, the sets are perfectly distinguishable by LOCC. But the consideration of hidden non-locality provides us the privilege to put a hierarchy among the sets. Precisely, we can claim that the sets of (5) are more local compared to those of (3), because the latter class contains hidden non-locality while for the former case there is no hidden non-locality though the set of (5) contains entangled states but the set of (3) does not.
@@ V Conclusion
In this manuscript, we have presented structures for locally distinguishable product states and entangled states such that they cannot be transformed to a locally indistinguishable set under orthogonality-preserving LOCC. Furthermore, we have constructed sets of multipartite states which are not activable in any bipartition. In other words, such locally distinguishable sets cannot be transformed to a locally indistinguishable set in any bipartition under orthogonality-preserving LOCC. This is, in fact, the worst case scenario in view of non-locality activation. We also have constructed locally distinguishable sets which can be transformed to locally indistinguishable sets under orthogonality-preserving LOCC. Then, we have classified the locally distinguishable states by introducing hierarchies. The structures that we have provided here can be easily generalized for high-dimensional Hilbert spaces. Finally, we have compared between locally distinguishable product states and entangled states.
@@ ACKNOWLEDGMENT
SH acknowledges partial funding by the European Union under Horizon Europe (grant agreement no. 101080086). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.